OpenStudy (anonymous):

find the domain of f(x)^-1 if f(x)=4x-8 ?

OpenStudy (mertsj):

f(x)=4x-8 is a line. Its inverse is also a line and thus the domain of both is all real numbers.

OpenStudy (anonymous):

thank you!!!

OpenStudy (anonymous):

No it isnt, be careful. 4x-8 is a line. But its inverse is \[{1 \over 4x-8}\] If the denominator becomes 0, the function explodes. You cant divide by zero. So in other words, 4x-8=0 may NOT happen. Solving for x: 4x = 8, x=2 So the domain is all the real numbers EXCEPT 2

OpenStudy (mertsj):

@soati \[y=4x-8\] Inverse is: \[x=4y-8\] \[x+8=4y\] \[y=\frac{1}{4}x+2=f ^{-1}(x)\]

OpenStudy (mertsj):

@soati A function and its inverse are symmetrical about the line y = x. Thus the inverse of a line is a line.

OpenStudy (mertsj):

@noniebug yw

OpenStudy (mertsj):

@soati You have written the reciprocal of the function, not the inverse. Perhaps you do not know the difference.

OpenStudy (anonymous):

lol really. I do. He wrote f(x)^-1 rather than f^-1(x). Thats a huge difference. \[f(x)^{-1} = 1/ f(x)\] \[f^{-1}(x)\] is what he wanted.

OpenStudy (mertsj):

@soati But what you wrote was that the INVERSE is \[\frac{1}{4x-8}\]

OpenStudy (anonymous):

He said f(x)^1, so i was correct. it doesnt matter what i said. the maths are correct. not sure why youre holding on to this.

OpenStudy (mertsj):

@soati Because you contradicted a true statement with a false one and you're still insisting that your false statement is true: " No it isnt, be careful. 4x-8 is a line. But its inverse is 1/ 4x−8" So who is holding on to what?

OpenStudy (anonymous):

holy crap. you either like the conflict or youre just retarded. wont be feeding your trolling anymore.